OpenStudy (anonymous):

A container holds a combination of nickels and quarters. There are 156 coins in all, and their total value is $21.60. The system of equations models this situation, where x is the number of nickels and y is the number of quarters. How many nickels and quarters are in the container? A. 118 nickels and 38 quarters B. 106 nickels and 50 quarters C. 87 nickels and 69 quarters D. 72 nickels and 84 quarters

OpenStudy (anonymous):

@BarfBag

OpenStudy (anonymous):

i dunno plz help if u dont know it is ok

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

It's basically using the system of equations. I will set you up with the main two equations to solve: \[\large x + y = 156\] \[\large 0.05x + 0.25y = $21.60\]

OpenStudy (anonymous):

Well then, there you have it!

OpenStudy (anonymous):

Hint: \[y=156-x\] Insert that into the other equation to solve.

OpenStudy (anonymous):

Basically, you want to render the equation 0.06 + 0.25y = $21.60 in either only the x-variable or the y-variable.

OpenStudy (anonymous):

0.05x*